OpenStudy (anonymous):

64 × 4^2 – 3 × 2^2=? help plz thx

OpenStudy (mertsj):

\[64\times4^2-3\times2^2\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[64\times16-3\times4\]

OpenStudy (anonymous):

PleaseExcuseMyDearAuntSally

OpenStudy (mertsj):

\[1024-12=1012\]

OpenStudy (campbell_st):

64 = 2^6 4 = 2^2 so 4^2 = 2^4 then 2^6 x 2^4 - 3 x 2^2 = 2^10 - 3^2

OpenStudy (campbell_st):

or 2^2( 2^9 - 3)

OpenStudy (anonymous):

=1012 right?

OpenStudy (anonymous):

1024 - 12 =1012